# Signal Processing for Haptic Surface Modeling: a Review

### v1

@@ 1 Introduction
The primary objective of extended-Reality (XR), in all its possible variants (virtual-, augmented-, mixed-) is to either faithfully replicate real-world experiences or enhance them [1, 2], enriching the real-world information by overlaying additional virtual contents[3, 4]. To effectively complement and/or replace their actual real-world counterparts, virtual experiences must be designed to engage as much as possible the various human senses [5].
@@
General [6, 7, 8, 9, 10, 11]
@@
(I) Sensors [12, 13, 14, 15]
@@
[16, 17, 18, 19, 20]
@@
[21, 22, 23, 24, 25, 26, 27, 28, 29]
@@
(V) Perception [30, 31, 32]
@@
Applications [33, 34, 35, 36, 37, 38, 39, 40]
@@
In this survey, we aim at reviewing such missing key aspect, i.e., the representation of real-world haptic features in the digital domain. We provide a comprehensive review on the representation of haptic surfaces from an image and signal processing perspective. We examine, which data are mostly used in today’s applications, looking at the existing datasets, and we indicate how they are being adopted by the research community. Drawing a parallelism with the research in the visual and acoustic domains, we highlight the shortcoming of the current haptic processing pipelines, pointing out possible research directions.
@@ 2.1 Sensing
Haptic sensors aim at mimicking the receptors that are present in our fingers and hands, to extract essential tactile information. Two main types of sensors exist: inertial and vision-based. Inertial sensors adopt various technologies (resistive, capacitive, fiber-based) to detect the roughness of the surface through the contact with it. They produce one-dimensional signals that capture the forces of interaction between sensor and object, which convey information about surface properties. Inertial sensors have been largely employed in robotic applications [13, 12], where tactile sensing refers to the point of contact between robot fingers and objects.
Vision-based sensors [15, 14] have recently emerged as the most simple and affordable way to capture tactile information at high resolution from images or video. The use of images and video appears to be very intuitive to tackle the problem. This is generally implemented using a RGB camera recording the deformation of a membrane upon its contact with a surface, thus generating a haptic map.
@@ 2.2 Haptic data processing and modeling
Haptic data processing and modeling, also called Haptic Surface Modeling, investigates how the sensed haptic signals are handled and stored in digital form. Unlike other sensory domains such as visual [41] and acoustics [42], where standardized representations are commonly available, the representation of tactile stimuli remains an area without established standards. In the early stages, haptic surface modeling relied on parametric functions to model the surface shape [43]. However, this approach either introduced a high amount of noise and approximation in the model or relied on too complex mathematical expressions. Consequently, researchers have opted for a paradigm change, moving towards data-driven approaches [44], where the object’s haptic information is collected as a one-dimensional signal, associated to classical representations such as spectrograms. The natural evolution of these initial approaches was then to turn to 2D signals, such as images and videos as well as haptic maps, which provide a much richer information that can be approached through well-studied computer vision techniques.
@@ 2.3 Representation and communication
Even more than its visual and audio counterparts, the haptic signal needs to rely on low latency communication to develop real-time interactive systems [19]. Potential applications include tele-surgery, industrial processes automation and augmented reality for live music concerts [16]. To enable such applications, researchers have focused on building the so-called Tactile Internet [16, 45, 19], elaborating on suitable technologies and communication protocols in the different layers of an architecture, such as the information representation using suitable codecs [45], as well as network design and customer’s haptic devices.
@@ 2.4 Haptic devices
Haptic devices aim at providing the tactile experience to the user, effectively translating digital data into physical stimuli that hit the receptors of human haptic channels [28]. The haptic stimulus can be delivered using single mode or multi-modal devices. Among single mode devices, surface haptic displays [25] consist of an active physical planar device that deforms so as to reproduce a surface model, usually sensed by the bare finger. Differently, multimodal haptic devices, like hand-held (joystick-like, unconstrained devices), wearables (usually gloves) and encountered/grounded (surfaces/devices for the user to touch ”on-demand”) [28, 22, 21, 29] aim at exploiting different stimuli, such as forces, vibrations and thermal, to convey the haptic sensation to the user.
@@ 2.5 Perception
Human haptic perception encompasses a range of sensory modalities, which can act externally in the form of cutaneous sensing, and internally as proprioception and kinesthesia [30]. While single mode haptic devices aim at reproducing pseudo-haptic feedback [31, 32], they struggle to fulfill the whole range of human haptic perception. Multi-modal devices could expand the spectrum of haptic feedback, reproducing the sensations produced by touching an object rather than its physical features, thus creating the so-called haptic illusions [30]. In human-centered applications, Quality of Experience (QoE) plays a crucial role in influencing the design and overall success of the whole pipeline. However, QoE in haptics is usually evaluated through subjective tests with the user-in-the-loop [45]. The desirable evaluation of QoE through objective testing aims at defining algorithmic models of human perception to calculate meaningful perceptual metrics [45, 7]. Other than human-based applications, perception is also key in robotics, for instance when dealing with object grasping [38, 46]. In such scenarios, understanding objects’ properties plays a fundamental role to modulate the applied force in the grasping action and to minimize failures.
@@ 3 Real world objects representation in the physical and the perceptual spaces
Fig. 2 illustrates the modeling pipeline, which connects physical phenomena sensed by specific sensors, the physical feature space that defines the object’s characteristics, and the perception feature space that identifies its properties as they are perceived by humans. While physical and perceptual spaces are well-established for images and audio signals, haptic signals still lack a standard representation able to characterize and replicate the relevant properties [47].
@@
Figure 2: The classic signal processing pipeline (top row) aims at mapping a physical phenomenon to a physical feature space such that a similar experience of the phenomenon can be delivered to the user in the perception feature space (bottom row). While the pipeline is well-established in both the visual and acoustic domains, it still lacks a common standard in the haptic domain.
@@ 3.1 Physical feature space
Visual and acoustic domains
The signals captured by cameras and microphones are mapped to standardised physical spaces. In the visual domain, light waves are filtered to obtain the RGB color model [41, 48], which resemble the wavelengths perceived by human eyes. In the acoustic domain, sound waves are one-dimensional signals, described by frequency, amplitude, and phase [42, 49].
@@
The properties describing the haptic qualities of an object in the physical space and a standard model encapsulating them did not reach a consensus among researchers yet. Nevertheless, as will be seen in the next paragraphs, researchers agree on the inclusion of some common macro categories as, for example, the concepts of roughness and friction.
The first attempt to define a haptic physical space and to correlate the physical properties of a material with their perception is described in [50]. In this work, the authors identified four physical properties that characterize a surface, namely: roughness, compliance, friction, and cooling rate. According to the authors, roughness measures the variations of the height of the surface, compliance measures its softness, friction measures the sliding resistance of the surface, and cooling rate measures the variations of temperature across it.
More recently, in [51, 52] the authors proposed a new way to characterize objects according to a set of haptic properties that can be considered orthogonal in what they called Haptic Attribute Space (HAS). In particular, the attributes that describe the four-dimensional HAS space are defined as follows: roughness (or micro-roughness) corresponds to the finer spaced irregularities of the surface texture that usually result from the inherent action of the production process or material condition [53]; bumpiness (also called macro-roughness or waviness) is associated to the macro-behavior of the surface texture, such as ridges, bumps and valleys; stiffness (or elasticity), it is the term used to describe the force needed to achieve a certain deformation of a structure [54]; friction (or slipperiness) refers to the resistance that one surface or object encounters when moving over another. In other words, it is the force that opposes the relative motion or tendency of such motion of two surfaces in contact [55]. In fig. 3, a visual example is depicted of the above properties.
Figure 3: In this example illustrating the section of a tree the red line is the roughness, the green line is the bumpiness, the blue line represents the stiffness, and the dark gray line is the friction.
@@ 3.2 Perception feature space
Visual and acoustic domains
The color feature space describing the human perceptual system is the CIELAB standard [56], which defines the relation between the wavelength distributions in the electromagnetic visible spectrum and the psychologically perceived colors in human vision [57]. In the acoustic domain, perception is characterized by pitch, loudness, and timbre, corresponding to frequency, amplitude, and phase in the physical space. Pitch relates to the perception of fundamental frequency [58, 49], loudness refers to sound intensity [59, 60], while timbre represents the tonal quality that distinguishes different sounds [61].
@@
Also in this case, differently from visual and acoustic stimuli, haptic stimulus is quite complex to describe from a perceptual viewpoint. In fact, understanding which are the surface properties within the physical space that influence our perception when interacting with an object is still an open problem [25]. Therefore, many studies have been conducted to identify an orthogonal perceptual feature space that optimizes the representation of the properties to transmit to our skin, allowing us to recognize and categorize the surfaces.
In early studies, such as [62], the authors suggested a possible two-dimensional space that describes the haptic perception in terms of roughness and hardness. Subsequently, slipperiness was added as a third dimension [63]. In [64], which has become the most common representation, the haptic features that compose the perceptual space were extended to four, namely roughness, compliance, coldness and friction. However, in [65], the authors further refined this model, proposing a 5-dimensional haptic perception space that included macro-roughness, micro-roughness, coldness, hardness, and friction. Other studies [66, 67] tried to apply PCA on a big set of properties classified by users, to identify the components that carry more information, in an attempt to minimise the dimensional space. In the opposite direction, other studies proposed to further enlarge the dimensional space. In [68], for instance, two new elements called ”volume” and ”naturalness”, were added to the perceptual space. According to the authors, ”volume” describes a shape property of the surface and ”naturalness” measures how much a surface is close to the natural world (as opposed to the artificial world). Another study analyzed the neural cortex response of primates and humans when scanning a surface [69]. The authors claim that the perceptual representation of a texture is high-dimensional and that the four properties usually adopted are not capable of conveying all the descriptors that are needed to uniquely describe a surface.
Notwithstanding all the recent research efforts, the four-dimensional model identified in 2010 [64] remains the one the research community still agrees more. Even when agreeing on these dimensions, however, it is still unclear which features carry more perceptual information (e.g., the L component in the CIELAB model conveys more information to the human eye than the other two), and how to model the relationship among features (e.g., in CIELAB, the non-linear relation between its components try to mimic the non-linear response of the human visual system).
@@ 3.3 Linking physical and perceptual feature spaces in haptics
While the haptic perception feature space has received broader attention from researchers, it is still unclear how to establish a direct link between the physical feature space and the perceptual feature space [70, 71]. A first attempt to define this mapping function was made in [71], where the authors worked at the mapping between textures acquired by a sensor and relevant features, such as roughness and bumpiness, as perceived by the users involved in their experiment.

### v2

@@ 1 Introduction
The primary objective of extended-Reality (XR), in all its possible variants (virtual-, augmented-, mixed-) is to either faithfully replicate real-world experiences or enhance them [xi2023challenges, alizadehsalehi2020bim], enriching the real-world information by overlaying additional virtual contents[keshner2004virtual, jungherr2022extended]. To effectively complement and/or replace their actual real-world counterparts, virtual experiences must be designed to engage as much as possible the various human senses [kral2022digital].
@@
General [biswas2021haptic, culbertson2018haptics, fleury2020survey, laycock2007survey, salisbury2004haptic, varalakshmi2012haptics]
@@
(I) Sensors [chen2018tactile, kappassov2015tactile, yamaguchi2019recent, zhang2022hardware]
@@
[promwongsa2020comprehensive, minopoulos2019survey, kokkonis2012survey, awais2023towards, antonakoglou2018toward]
@@
[adilkhanov2022haptic, basdogan2020review, breitschaft2022haptic, choi2012vibrotactile, costes2020towards, mercado2021haptics, tan2020methodology, wang2019multimodal, wee2021haptic]
@@
(V) Perception [kurzweg2024survey, lecuyer2009simulating, ujitoko2021survey]
@@
Applications [coles2010role, escobar2016review, gaffary2018use, hamza2019haptic, imran2021significance, luo2017robotic, petermeijer2015effect, saint2021survey]
@@
In this work we focus on the problem of associating the visual aspect of a textured surface (its image) to the relevant haptic information and we review the signal processing methodologies that allow to extract and model such information. This is a fundamental aspect for the image/video communication systems of the future and for the construction of truly immersive AR/VR systems [chen2024comprehensive]. Indeed, although some attempts have been proposed in the past [kim2013construction], at present there are no standards for managing haptic information in multimedia content streaming. We believe that the progress in this field builds upon consolidated and universally accepted methodologies to associate image and haptic information, and signal processing is certainly a key enabling technology in this context.
In this survey, we aim at reviewing such missing key aspect, i.e., the representation of real-world haptic features in the digital domain. We provide a comprehensive review on the representation of haptic surfaces from an image and signal processing perspective. We examine, which data are mostly used in today’s applications, looking at the existing datasets, and we indicate how they are being adopted by the research community. Drawing a parallelism with the research in the visual and acoustic domains, we highlight the shortcoming of the current haptic processing pipelines, pointing out possible research directions. Although the sense of touch can be experienced across the entire human body [ke2022propelwalker, ushiyama2023feetthrough], this survey focuses primarily on hand interactions, which are prevalent in the available literature and commercial devices.
@@ 2.1 Sensing
Haptic sensors aim at mimicking the receptors that are present in our fingers and hands, to extract essential tactile information. Although many types of sensors exist to collect haptic data, the two types of sensors majorly employed to model haptic properties are inertial and vision-based. Inertial sensors adopt various technologies (resistive, capacitive, fiber-based) to detect the roughness of the surface through the contact with it. They produce one-dimensional signals that capture the forces of interaction between sensor and object, which convey information about surface properties. Inertial sensors have been largely employed in robotic applications [kappassov2015tactile, chen2018tactile], where tactile sensing refers to the point of contact between robot fingers and objects.
Vision-based sensors [yamaguchi2019recent, zhang2022hardware] have recently emerged as the most simple and affordable way to capture tactile information at high resolution from images or video. The use of images and video appears to be very intuitive to tackle the problem. This is generally implemented using a RGB camera recording the deformation of a membrane upon its contact with a surface, thus generating a haptic map.
@@ 2.2 Haptic data processing and modeling
Haptic data processing and modeling, also called Haptic Surface Modeling, investigates how the sensed haptic signals are handled and stored in digital form. Unlike other sensory domains such as visual [susstrunk1999standard] and acoustics [fletcher2012physics], where standardized representations are commonly available, the representation of tactile stimuli remains an area without established standards. In the early stages, haptic surface modeling relied on parametric functions to model the surface shape [basdogan2002haptic]. However, this approach either introduced a high amount of noise and approximation in the model or relied on too complex mathematical expressions. Consequently, researchers have opted for a paradigm change, moving towards data-driven approaches [hover2008data], where the object’s haptic information is collected as a one-dimensional signal, associated to classical representations such as spectrograms. The natural evolution of these initial approaches was then to turn to 2D signals, such as images and videos as well as haptic maps, which provide a much richer information that can be approached through well-studied computer vision techniques.
@@ 2.3 Representation and communication
The representation of haptic data is still underexplored, complicating the integration of haptic stimuli within virtual environments. Although beyond the scope of this survey, we provide a brief overview of the literature addressing the representation of haptic information. In [wakita2008texturebased], the authors define a method to encode haptic information into 3 maps, each representing a specific haptic stimulus. These maps include height maps, friction maps, and stiffness maps [wakita2008texturebased, wakita2010friction]. Similarly to this approach, in [kamuro2012haptic], the authors developed an editor that uses four layers to define the haptic stimuli. In [kim2011haptic], the authors extract haptic information from visual images and represent it using depth, stiffness, and viscosity maps.
In [costes2018haptic], the authors draw a more complete overview of the previous methods. Furthermore, they define 10 maps correlating haptic characteristics (e.g., compliance, geometry) with perceptual modalities (e.g., kinesthetic, thermal).
Even more than its visual and audio counterparts, the haptic signal needs to rely on low latency communication to develop real-time interactive systems [awais2023towards]. Potential applications include tele-surgery, industrial processes automation and augmented reality for live music concerts [promwongsa2020comprehensive]. To enable such applications, researchers have focused on building the so-called Tactile Internet [promwongsa2020comprehensive, steinbach2018haptic, awais2023towards], elaborating on suitable technologies and communication protocols in the different layers of an architecture, such as the information representation using suitable codecs [steinbach2018haptic], as well as network design and customer’s haptic devices.
@@ 2.4 Haptic devices
Haptic devices aim at providing the tactile experience to the user, effectively translating digital data into physical stimuli that hit the receptors of human haptic channels [wang2019multimodal]. The haptic stimulus can be delivered using single mode or multi-modal devices. Among single mode devices, surface haptic displays [costes2020towards] consist of an active physical planar device that deforms so as to reproduce a surface model, usually sensed by the bare finger. Differently, multimodal haptic devices, like hand-held (joystick-like, unconstrained devices), wearables (usually gloves) and encountered/grounded (surfaces/devices for the user to touch ”on-demand”) [wang2019multimodal, basdogan2020review, adilkhanov2022haptic, wee2021haptic] aim at exploiting different stimuli, such as forces, vibrations and thermal, to convey the haptic sensation to the user.
@@ 2.5 Perception
Human haptic perception encompasses a range of sensory modalities, which can act externally in the form of cutaneous sensing, and internally as proprioception and kinesthesia [kurzweg2024survey]. While single mode haptic devices aim at reproducing pseudo-haptic feedback [lecuyer2009simulating, ujitoko2021survey], they struggle to fulfill the whole range of human haptic perception. Multi-modal devices could expand the spectrum of haptic feedback, reproducing the sensations produced by touching an object rather than its physical features, thus creating the so-called haptic illusions [kurzweg2024survey]. In human-centered applications, Quality of Experience (QoE) plays a crucial role in influencing the design and overall success of the whole pipeline. However, QoE in haptics is usually evaluated through subjective tests with the user-in-the-loop [steinbach2018haptic]. The desirable evaluation of QoE through objective testing aims at defining algorithmic models of human perception to calculate meaningful perceptual metrics [steinbach2018haptic, culbertson2018haptics]. Other than human-based applications, perception is also key in robotics, for instance when dealing with object grasping [luo2017robotic, li2020review]. In such scenarios, understanding objects’ properties plays a fundamental role to modulate the applied force in the grasping action and to minimize failures.
@@ 3 Real world objects representation in the physical and the perceptual spaces
Fig. 2 illustrates the modeling pipeline, which connects physical phenomena sensed by specific sensors, the physical feature space that defines the object’s characteristics, and the perception feature space that identifies its properties as they are perceived by humans. While physical and perceptual spaces are well-established for images and audio signals, haptic signals still lack a standard representation able to characterize and replicate the relevant properties [degraen2021capturing].
@@
Figure 2: The classic signal processing pipeline (top row) aims at mapping a physical phenomenon to a physical feature space such that a similar experience of the phenomenon can be delivered to the user in the perception feature space (bottom row). While the pipeline is well-established in both the visual and acoustic domains, it still lacks a common standard in the haptic domain. In the example in bottom row illustrating the section of a tree, the red line is the roughness, the green line is the bumpiness, the blue line represents the stiffness, and the dark gray line is the friction.
@@ 3.1 Physical feature space
Visual domain
The RGB color model [susstrunk1999standard] is the most widespread model to describe the physical lights properties in the visual domain [susstrunk1999standard, ibraheem2012understanding]. This model identifies three main wavelengths perceived by our eye, which are then combined to synthesize all the other colors of the visible spectrum. To convert colors to a digital representation, a digital RGB camera measures the quantity of photons hitting the sensor. Then, the sensors converts the photons quantity into an electric current which voltage value is quantized to obtain to the digital representation of the original color.
Acoustic domain
Acoustic waves are the signal generated by the vibrations of a source. Their physical characteristics are usually described by 3 quantities known as frequency, amplitude and phase [fletcher2012physics]. More specifically, the frequency describes the pitch of the sound, the amplitude describes its strength, and the phase describes the position of a sound wave relative to a reference point in its cycle [fletcher2012physics, muller2015fundamentals]. Similarly to the visual domain, the acoustic waves are first collected by a microphone equipped with a membrane, whose vibrations are converted to an electric signal and then digitized.
@@
The properties describing the haptic qualities of an object in the physical space and a standard model encapsulating them did not reach a consensus among researchers yet. Nevertheless, as will be seen in the next paragraphs, researchers agree on the inclusion of some common macro categories as, for example, the concepts of roughness and friction. In fact, researchers have focused more on the perception feature space, living behind the studies aiming at characterizing which are the physical properties that can be used to describe the haptic sensations.
The first attempt to define a haptic physical space and to correlate the physical properties of a material with their perception is described in [chen2009exploring]. In this work, the authors identified four physical properties that characterize a surface, namely: roughness, compliance, friction, and cooling rate. According to the authors, roughness measures the variations of the height of the surface, compliance measures its softness, friction measures the sliding resistance of the surface, and cooling rate measures the variations of temperature across it.
More recently, in [hassan2019authoring, hassan2023establishing] the authors proposed a new way to characterize objects according to a set of haptic properties that can be considered orthogonal in what they called Haptic Attribute Space (HAS). In particular, the attributes that describe the four-dimensional HAS space are defined as follows: roughness (or micro-roughness) corresponds to the finer spaced irregularities of the surface texture that usually result from the inherent action of the production process or material condition [american2003surface]; bumpiness (also called macro-roughness or waviness) is associated to the macro-behavior of the surface texture, such as ridges, bumps and valleys; stiffness (or elasticity), it is the term used to describe the force needed to achieve a certain deformation of a structure [baumgart2000stiffness]; friction (or slipperiness) refers to the resistance that one surface or object encounters when moving over another. In other words, it is the force that opposes the relative motion or tendency of such motion of two surfaces in contact [bowden2001friction]. In Fig. 2, a visual example is depicted of the above properties.
@@ 3.2 Perception feature space
Visual domain
The CIELAB standard [standard2007colorimetry, cheung2012uniform], an evolution of the initially proposed CIE XYZ 1931 [smith1931colorimetric], aims at quantitatively describing the relation between the wavelength distributions in the electromagnetic visible spectrum and the physiologically perceived colors in human vision [stiles1959npl]. The CIELAB color space is built starting from the human vision and, differently from the RGB space, it is device-independent (e.g., the same RGB triplet can look different on separate devices, and it is thus mapped differently in the CIELAB space), although almost all devices currently rely on automatic gamut adaptation to the device (like sRGB). The relation between RGB and CIELAB spaces has been defined in [connolly1997study]. Subsequent works [fdhal2009color, ibraheem2012understanding] have continued the investigation on the relationship between the physical and perception spaces as well as the classification of the color spaces and how they are built.
Acoustic domain
The perception space in the acoustic domain is characterised by 3 main features: Pitch, Loudness, and Timbre [muller2015fundamentals, moore2012introduction]. Differently from the acoustic physical space, these features have been already used to define a three-dimensional space, where sound waves can be considered as vectors. However, many studies correlated these properties to the physical space dimensions, namely: Frequency, Amplitude and Phase. The Pitch of a sound is closely related to the frequency notion, but it refers to our perception of the fundamental frequency [zwicker2013psychoacoustics, muller2015fundamentals]. The Loudness dimension refers to the sound intensity [fletcher1933loudness, moore2012introduction] and its value varies on a scale from quiet to loud [muller2015fundamentals]. The Timbre [sethares2005tuning] is the characteristic allowing us to recognize the tone color of the sound we are hearing.
@@
Also in this case, differently from visual and acoustic stimuli, haptic stimulus is quite complex to describe from a perceptual viewpoint. In fact, understanding which are the surface properties within the physical space that influence our perception when interacting with an object is still an open problem [costes2020towards]. Therefore, many studies have been conducted to identify an orthogonal perceptual feature space that optimizes the representation of the properties to transmit to our skin, allowing us to recognize and categorize the surfaces.
In early studies, such as [holliins1993perceptual], the authors suggested a possible two-dimensional space that describes the haptic perception in terms of roughness and hardness. Subsequently, slipperiness was added as a third dimension [hollins2000individual]. In [tiest2010tactual], which has become the most common representation, the haptic features that compose the perceptual space were extended to four, namely roughness, compliance, coldness and friction. However, in [okamoto2012psychophysical], the authors further refined this model, proposing a 5-dimensional haptic perception space that included macro-roughness, micro-roughness, coldness, hardness, and friction. Other studies [baumgartner2013visual, metzger2022unsupervised] tried to apply PCA on a big set of properties classified by users, to identify the components that carry more information, in an attempt to minimise the dimensional space. In the opposite direction, other studies proposed to further enlarge the dimensional space. In [sakamoto2017exploring], for instance, two new elements called ”volume” and ”naturalness”, were added to the perceptual space. According to the authors, ”volume” describes a shape property of the surface and ”naturalness” measures how much a surface is close to the natural world (as opposed to the artificial world). Another study analyzed the neural cortex response of primates and humans when scanning a surface [lieber2019high]. The authors claim that the perceptual representation of a texture is high-dimensional and that the four properties usually adopted are not capable of conveying all the descriptors that are needed to uniquely describe a surface.
Notwithstanding all the recent research efforts, the four-dimensional model identified in 2010 [tiest2010tactual] remains the one the research community still agrees more. Even when agreeing on these dimensions, however, it is still unclear which features carry more perceptual information (e.g., the L component in the CIELAB model conveys more information to the human eye than the other two), and how to model the relationship among features (e.g., in CIELAB, the non-linear relation between its components try to mimic the non-linear response of the human visual system).
@@ 3.3 Linking physical and perceptual feature spaces in haptics
While the haptic perception feature space has received broader attention from researchers, it is still unclear how to establish a direct link between the physical feature space and the perceptual feature space [lu2022preference, awan2023predicting]. A first attempt to define this mapping function was made in [awan2023predicting], where the authors worked at the mapping between textures acquired by a sensor and relevant features, such as roughness and bumpiness, as perceived by the users involved in their experiment.